# GeoZero: Incentivizing Reasoning from Scratch on Geospatial Scenes

### v1

@@ 1 Introduction
Geospatial data, represented primarily by remote sensing imagery, provides a comprehensive visual perspective on the Earth’s surface. However, interpreting such data remains challenging due to the diverse scales, shapes, and spatial layouts of geospatial objects [Wang_2025_CVPR_XLRSBench, gong2024crossearth], as well as variations in illumination, sensor characteristics, and environmental conditions [an2025choice, wang2025hypersigma].
Multimodal large language models (MLLMs) [liu2023visual, zhu2023minigpt, alayrac2022flamingo, Dai2023InstructBLIPTG, bai2023qwen] have been introduced to enhance geospatial understanding through instruction tuning on remote sensing datasets [kuckreja2024geochat, pang2025vhm, luo2024skysensegpt]. Although these models can handle basic remote sensing dialogues, they often fail in complex or unseen environments. This limitation mainly stems from relying solely on supervised fine-tuning (SFT), which leads the model to memorize linguistic patterns rather than develop genuine geospatial reasoning [sft_memory_rl_generalize].
@@
Recently, advanced language models such as GPT-o1 [gpt-o1] and DeepSeek-R1 [deepseek-r1] have demonstrated explicit reasoning capabilities, characterized by performing deliberate thought before producing answers. Inspired by this progress, the remote sensing community has begun to explore MLLMs capable of “thinking” about geospatial scenes [koksal2025tinyrs, scorers-r1]. Most existing approaches follow the DeepSeek-R1 training paradigm (see Figure 1): (1) constructing chain-of-thought (CoT) data for SFT to initialize the model with a basic “thinking–answering” response pattern, and (2) applying reinforcement learning (RL), such as Group Relative Policy Optimization (GRPO), to further enhance its reasoning capability. In these frameworks, the CoT annotations are typically generated by advanced general-purpose MLLMs (e.g., GPT-4o [hurst2024gpt4o]) and subsequently verified or refined through manual inspection.
By examining the above process, we argue that, beyond the substantial annotation cost arising from the professional nature of remote sensing interpretation [dotav2], training on manually revised reasoning paths inherently limits the model’s cognitive potential. In our view, manually refining CoTs inevitably introduces subjective biases that may constrain the model’s ultimate performance [humanbias_emnlp, llm_bias, human_ai_act_bias]. To this end, a critical question naturally arises:
@@
To address this question, we present GeoZero, a MLLM capable of performing entirely internal reasoning on geospatial scenes without relying on any predefined CoT supervision. To achieve this, we follow the standard SFT–RL training paradigm commonly used by multimodal reasoning models [koksal2025tinyrs, vision-r1] and construct two remote sensing vision–language datasets: GeoZero-Instruct and GeoZero-Hard (both containing only question–answer pairs without any CoT data). In the SFT stage, GeoZero-Instruct adapts general-purpose MLLMs to geospatial scenarios, enabling the acquisition of fundamental domain knowledge. In the RL stage, since simple questions can often be answered without genuine reasoning, GeoZero-Hard is employed to promote deep reasoning by exposing the model to the most challenging examples. Furthermore, we introduce an enhanced RL optimization framework, Answer-Anchored GRPO (A2GRPO), where “anchored” denotes the joint refinement of reasoning quality and answer correctness based on the model’s predicted answers, thereby encouraging diverse yet accurate reasoning.
@@
(3) We conduct extensive experiments on various remote sensing vision–language benchmarks, where GeoZero not only achieves competitive performance compared with existing methods but also exhibits emergent deep reasoning. Before answering, the model generates reasoning trajectories entirely derived from its own internal cognition, demonstrating adaptive reasoning capabilities across diverse geospatial tasks.
@@ 2.1 MLLMs on Geospatial Scenes
Benefiting from the rapid progress of general-purpose LLMs and MLLMs [gpt3, touvron2023llama, liu2024deepseek, Dai2023InstructBLIPTG, liu2023visual, zhu2023minigpt], numerous geospatial-oriented MLLMs have recently been proposed [hu2025rsgpt, kuckreja2024geochat, zhan2025skyeyegpt, liu2024rsunivlm, wang2024ringmogpt, wang2025geollavak, soni2024earthdial]. Due to the scarcity of textual annotations in the remote sensing community, early efforts relied on manual captioning [hu2025rsgpt] or label-to-text conversion strategies [zhang2024earthgpt] to construct image–text pairs for model training. Unlike these manually generated datasets, later methods employed frontier LLMs or MLLMs to perform automatic textual annotation [kuckreja2024geochat, li2025lhrs, pang2025vhm, luo2024skysensegpt], significantly improving data generation efficiency. Recent works have also addressed domain-specific challenges such as large-size [wang2025geollavak, lrsvqa], multi-temporal [irvin2025teochat, li2024unirs], and multi-sensor [soni2024earthdial, shu2025earthmind] imagery. Beyond image–text dialogue systems, several MLLMs have been developed for vision-centric interpretation tasks [zhou2024geoground, si2025spex, shabbir2025geopixel, ou2025geopix, zheng2025instructsam]. However, none of these models are capable of generating explicit reasoning chains before producing an answer. Their knowledge primarily stems from memorized textual patterns acquired through SFT, which limits their ability to generalize to complex and unseen geospatial scenarios.
@@ 2.2 Large Multimodal Reasoning Models
The reasoning capability of large language models (LLMs) [gpt-o1, deepseek-r1] has been greatly enhanced through post-training [post-training], achieving impressive performance on complex logical tasks such as mathematical problem solving [yang2025qwen3] and science exploration [llmchemistry]. Especially, encouraged by the success of DeepSeek-R1 [deepseek-r1], numerous approaches have adopted GRPO to enhance the reasoning capability of MLLMs [vision-r1, peng2025lmm, mm-eureka, visual-rft, vlm-r1, r1-vl]. In the remote sensing domain, several preliminary explorations have also emerged [koksal2025tinyrs, geo-r1, geovlm-r1, scorers-r1, rsthinker, segearth-r1, remotereasoner, uav-vl-r1]. TinyRS-R1 [koksal2025tinyrs] conducts GRPO on models fine-tuned with its constructed reasoning dataset, where the CoT is generated by GPT-4.1-mini through prompting with images, question–answer pairs, and system messages. ScoreRS [scorers-r1] and Geo-R1 [geo-r1] further improve reasoning performance on visual grounding and few-shot referring expression tasks. Compared with the above methods, our model is capable of performing emergent multimodal reasoning across multiple remote sensing vision–language tasks. It generates thinking trajectories entirely from its own internal cognition, without relying on any CoT annotations, thereby enabling emergent reasoning in geospatial understanding.
@@ 3.1 Overall Training Paradigm
Figure 2 illustrates the overall framework of GeoZero, which employs a standard SFT–RL training paradigm. Following common practices in remote-sensing MLLMs [zhan2025skyeyegpt, zhang2024earthgpt], we first conduct an SFT without CoT annotations on a constructed large-scale instruction-tuning dataset, GeoZero-Instruct, to equip the model with a preliminary understanding of geospatial scenes. In the subsequent RL stage, we construct a dedicated hard-sample pool, GeoZero-Hard, exposing the model to more challenging cases to stimulate deep reasoning. This two-stage framework allows the model to first acquire general geospatial knowledge and then stimulate its reasoning ability on difficult examples.
@@ 3.2 Dataset Construction
To establish the foundation for model training, we first aggregate a wide range of publicly available remote sensing datasets into a unified collection, referred to as GeoZero-Raw. Based on this collection, we further derive two specialized subsets: GeoZero-Instruct for SFT and GeoZero-Hard for RL.
@@ 3.2.1 GeoZero-Raw
We mainly focus on four representative remote sensing vision–language tasks: scene classification (SC), visual grounding (VG), visual question answering (VQA), and image captioning (IC). For each task, we collect samples from multiple classical datasets. For instance, the SC task includes data from the training sets of AID [aid], RESISC-45 [nwpu], and NASC-TG2 [nasc], among others. The resulting GeoZero-Raw dataset spans diverse resolutions, sensor modalities, and scene categories, which helps enhance the generalization ability of MLLMs for remote sensing understanding. A detailed composition of the dataset is summarized in Table 1.
Instruction Formatting. After dataset aggregation, each sample is converted into an instruction-following format. The details can be found in the supplementary material.
It should be noted that during the dataset aggregation process, some samples may share the same imagery as those from other datasets. However, since their textual instructions are regenerated according to our instruction conversion rules (detailed in the supplementary material), these samples are treated as distinct entries. In total, the resulting GeoZero-Raw dataset contains about 750K samples.
@@ 3.2.2 GeoZero-Instruct and GeoZero-Hard
We design a three-step process to build these datasets: (1) train the MLLM on GeoZero-Raw to obtain a data filtering model (DFM); (2) select difficult samples to construct GeoZero-Hard; and (3) perform data filtering based on GeoZero-Hard to produce GeoZero-Instruct. Intuitively, according to the maximum likelihood principle, samples mispredicted in a single inference pass are more likely to be difficult. Therefore, to ensure both the difficulty of GeoZero-Hard and the efficiency of sample selection, we adopt a coarse-to-fine two-stage filtering strategy. First, all samples in GeoZero-Raw are fed into the DFM for one inference pass, and only those incorrectly predicted are retained as potential hard-sample candidates. Next, these candidates are re-evaluated three times using the same model with different random seeds. Finally, we rank all candidates by their average error rate and select the top portion as the final hard samples. To maintain task balance, an equal number of samples are drawn from each of the four primary tasks. The selected samples collectively constitute the hard-sample pool, referred to as GeoZero-Hard.
Notably, to avoid potential image overlap within GeoZero-Raw (as discussed in Section 3.2.1), we remove all samples sharing the same image with GeoZero-Hard. The remaining samples form GeoZero-Instruct. As a result, GeoZero-Hard and GeoZero-Instruct contain approximately 20K and 610K samples, respectively.
@@ 3.3 Answer-Anchored Group Relative Policy Optimization
Task Type Dataset #Samples Subtotal
Scene Classification VHM-Instruct [pang2025vhm] 9,050 110,649
AID-train [aid] 8,000
RESISC45-train [nwpu] 25,200
NASC-TG2-train [nasc] 16,000
WHU-RS19-train [whurs19] 799
EuroSAT-train [eurosat] 21,600
fMoW-RGB-train [fmow] 30,000
Visual Grounding VHM-Instruct [pang2025vhm] 53,982 126,620
RSVG-train [rsvg] 5,505
DIOR-RSVG-train [dior-rsvg] 30,820
VRS-train [li2024vrsbench] 36,313
Visual Question Answering VHM-Instruct [pang2025vhm] 10,000 243,041
RSVQA-HR-train [rsvqa] 80,000
RSVQA-LR-train [rsvqa] 67,228
VRS-train [li2024vrsbench] 85,813
Image Captioning SkyEye968k [zhan2025skyeyegpt] 153,375 173,639
VRS-train [li2024vrsbench] 20,264
Others VHM-Instruct [pang2025vhm] 100,800 100,800
Total - - 754,749
@@ 3.3.1 Reward Function
Answer-Modulated Thinking Reward. AMTR begins with a thinking-quality score $s_{t}$, which comprehensively evaluates multiple aspects of the reasoning process, including trajectory length, semantic diversity, sentence redundancy, and the degree of overlap with the final answer. The supplementary material provides more details of $s_{t}$.
@@
Task-Specific Answer Reward. To adapt MLLMs to remote sensing vision-language tasks, following existing practices [koksal2025tinyrs, geovlm-r1, uav-vl-r1], we additionally design task-specific answer rewards $r_{a}$ for SC, VG, VQA, and IC. For consistency in reward computation, $r_{a}$ is expected to fall within the range of $[0,1]$. Moreover, $r_{a}$ should preferably be continuous so that it can smoothly reflect the quality differences among various answers, thereby facilitating model optimization during RL training. To this end, we formulate dedicated reward functions for each task, and the detailed formulations are provided in the supplementary material.
@@ 3.3.2 Optimization Objective
In the original GRPO framework [deepseek-r1], the optimization objective depends on the logits of responses containing both the reasoning trajectory and the final answer. However, since our goal is to incentivize MLLMs to think from scratch, i.e., without exposure to any CoT data, the reference model (the MLLM after SFT) does not explicitly generate reasoning text. In this case, directly applying GRPO’s original formulation may inadvertently constrain the reasoning process. To address this issue, we refine the optimization objective in A2GRPO as follows:
@@
The definitions of $G$, $\varepsilon$, $\beta$, and the advantage term $A_{i}(r_{1},\dots,r_{G})$ follow the original GRPO [deepseek-r1], where $r_{i}$ denotes the overall reward (see Equation (3)) for each candidate response. The key difference lies in the introduction of a binary mask $M_{i}=\{m_{i,t}\}_{t=1}^{|o_{i}|}$, where $m_{i,t}=1$ if token $t$ is located within the answer tags (as detailed later), and $m_{i,t}=0$ otherwise. This design ensures that the distributional alignment with the reference model is enforced only on the answer portion, thereby preventing the reasoning tokens from being overly regularized and encouraging MLLMs to freely explore diverse reasoning trajectories.
@@ 3.3.3 Prompt template
Despite these efforts, MLLMs may still fail to generate explicit reasoning texts. In our experiments, we observed that directly using the original GRPO template [deepseek-r1] (<think>reasoning process</think><answer>answer</answer>) often fails to activate reasoning behaviors, as tokens ”<think>” or ”<” rarely appear at the beginning of sentences for SFT. Consequently, the model struggles to initiate coherent reasoning from scratch. To alleviate this issue, we simplify the GRPO prompt template by removing corresponding constraints while retaining only the ”<answer></answer>” markers. More details of our system prompts are provided in the supplementary material.

### v2

@@ 1 Introduction
Geospatial data, represented primarily by remote sensing imagery, provides a comprehensive visual perspective on the Earth’s surface. However, interpreting such data remains challenging due to the diverse scales, shapes, and spatial layouts of geospatial objects [59, 16], as well as variations in illumination, sensor characteristics, and environmental conditions [2, 57].
Multimodal large language models (MLLMs) [35, 76, 1, 11, 3] have been introduced to enhance geospatial understanding through instruction tuning on remote sensing datasets [28, 47, 41]. Although these models can handle basic remote sensing dialogues, they often fail in complex or unseen environments. This limitation mainly stems from relying solely on supervised fine-tuning (SFT), which leads the model to memorize linguistic patterns rather than develop genuine geospatial reasoning [10].
@@
Recently, advanced language models such as GPT-o1 [26] and DeepSeek-R1 [18] have demonstrated explicit reasoning capabilities, characterized by performing deliberate thought before producing answers. Inspired by this progress, the remote sensing community has begun to explore MLLMs capable of “thinking” about geospatial scenes [27, 45]. Most existing approaches follow the DeepSeek-R1 training paradigm (see Figure 1): (1) constructing chain-of-thought (CoT) data for SFT to initialize the model with a basic “thinking–answering” response pattern, and (2) applying reinforcement learning (RL), such as Group Relative Policy Optimization (GRPO), to further enhance its reasoning capability. In these frameworks, the CoT annotations are typically generated by advanced general-purpose MLLMs (e.g., GPT-4o [24]) and subsequently verified or refined through manual inspection.
By examining the above process, we argue that, beyond the substantial annotation cost arising from the professional nature of remote sensing interpretation [13], training on manually revised reasoning paths inherently limits the model’s cognitive potential. In our view, manually refining CoTs inevitably introduces subjective biases that may constrain the model’s ultimate performance [7, 43, 15]. To this end, a critical question naturally arises:
@@
A preliminary analysis (see Section-B in the supplementary material) shows that general instruction-tuned MLLMs exhibit extremely limited geospatial reasoning ability. To address this question, we present GeoZero, a MLLM capable of performing entirely internal reasoning on geospatial scenes without relying on any predefined CoT supervision. To achieve this, we follow the standard SFT–RL training paradigm commonly used by multimodal reasoning models [27, 23] and construct two remote sensing vision–language datasets: GeoZero-Instruct and GeoZero-Hard (both containing only question–answer pairs without any CoT data). In the SFT stage, GeoZero-Instruct adapts general-purpose MLLMs to geospatial scenarios, enabling the acquisition of fundamental domain knowledge. In the RL stage, since simple questions can often be answered without genuine reasoning, GeoZero-Hard is employed to promote deep reasoning by exposing the model to the most challenging examples. Furthermore, we introduce an enhanced RL optimization framework, Answer-Anchored GRPO (A2GRPO), where “anchored” denotes the joint refinement of reasoning quality and answer correctness based on the model’s predicted answers, thereby encouraging diverse yet accurate reasoning.
@@
(3) We conduct extensive experiments on various remote sensing vision–language benchmarks, where GeoZero not only achieves competitive performance compared with existing methods but also exhibits emergent deep reasoning. Before answering, GeoZero generates reasoning trajectories from its internal cognition, demonstrating adaptive reasoning capabilities across diverse geospatial tasks.
@@ 2.1 MLLMs on Geospatial Scenes
Benefiting from the rapid progress of general-purpose MLLMs [6, 56, 34, 11, 35, 76], numerous geospatial-oriented MLLMs have recently been proposed [22, 28, 68, 37, 60, 58, 54]. Due to the scarcity of textual annotations in the remote sensing community, early efforts relied on manual captioning [22] or label-to-text conversion strategies [70] to construct image–text pairs for model training. Unlike these manually generated datasets, later methods employed frontier MLLMs to perform automatic textual annotation [28, 33, 47, 41], significantly improving data generation efficiency. Recent works have also addressed domain-specific challenges such as large-size [58, 42], multi-temporal [25, 32], and multi-sensor [54, 52] imagery. Beyond image–text dialogue systems, several MLLMs have been developed for vision-centric interpretation tasks [74, 53, 50, 46, 73]. However, none of these models are capable of generating explicit reasoning chains before producing an answer. Their knowledge primarily stems from memorized textual patterns acquired through SFT, which limits their ability to generalize to complex and unseen geospatial scenarios.
@@ 2.2 Large Multimodal Reasoning Models
The reasoning capability of large language models (LLMs) [26, 18] has been greatly enhanced through post-training [29], achieving impressive performance on complex logical tasks such as mathematical problem solving [64] and science exploration [71]. Especially, encouraged by the success of DeepSeek-R1 [18], numerous approaches have adopted GRPO to enhance the reasoning capability of MLLMs [23, 48, 44, 38, 51, 69]. In the remote sensing domain, several preliminary explorations have also emerged [27, 72, 14, 45, 36, 30, 66, 17]. TinyRS-R1 [27] conducts GRPO on models fine-tuned with its constructed reasoning dataset, where the CoT is generated by GPT-4.1-mini through prompting with images, question–answer pairs, and system messages. ScoreRS [45] and Geo-R1 [72] further improve reasoning performance on visual grounding and few-shot referring expression tasks. Compared with the above methods, our model is capable of performing emergent multimodal reasoning across multiple remote sensing vision–language tasks. It generates thinking trajectories entirely from its own internal cognition, without relying on any CoT annotations, thereby enabling emergent reasoning in geospatial understanding.
@@ 3.1 Overall Training Paradigm
Figure 2 illustrates the overall framework of GeoZero, which employs a standard SFT–RL training paradigm. Following common practices in remote-sensing MLLMs [68, 70], we first conduct an SFT without CoT annotations on a constructed large-scale instruction-tuning dataset, GeoZero-Instruct, to equip the model with a preliminary understanding of geospatial scenes. In the subsequent RL stage, we construct a dedicated hard-sample pool, GeoZero-Hard, exposing the model to more challenging cases to stimulate deep reasoning. This two-stage framework allows the model to first acquire general geospatial knowledge and then stimulate its reasoning ability on difficult examples.
@@
Task Type Dataset #Samples Subtotal
Scene Classification VHM-Instruct [47] 9,050 110,649
AID-train [61] 8,000
RESISC45-train [8] 25,200
NASC-TG2-train [75] 16,000
WHU-RS19-train [62] 799
EuroSAT-train [19] 21,600
fMoW-RGB-train [9] 30,000
Visual Grounding VHM-Instruct [47] 53,982 126,620
RSVG-train [55] 5,505
DIOR-RSVG-train [67] 30,820
VRS-train [31] 36,313
Visual Question Answering VHM-Instruct [47] 10,000 243,041
RSVQA-HR-train [39] 80,000
RSVQA-LR-train [39] 67,228
VRS-train [31] 85,813
Image Captioning SkyEye968k [68] 153,375 173,639
VRS-train [31] 20,264
Others VHM-Instruct [47] 100,800 100,800
Total - - 754,749
@@ 3.2 Dataset Construction
To establish the foundation for model training, we first aggregate publicly available remote sensing datasets into a unified collection, referred to as GeoZero-Raw. Based on this collection, we further derive two specialized subsets: GeoZero-Instruct for SFT and GeoZero-Hard for RL.
@@ 3.2.1 GeoZero-Raw
We mainly focus on four representative remote sensing vision–language tasks: scene classification (SC), visual grounding (VG), visual question answering (VQA), and image captioning (IC). For each task, we collect samples from multiple classical datasets. For instance, the SC task includes data from the training sets of AID [61], RESISC-45 [8], and NASC-TG2 [75], among others. The resulting GeoZero-Raw dataset spans diverse resolutions, sensor modalities, and scene categories, which helps enhance the generalization ability of MLLMs for remote sensing understanding. A detailed composition of the dataset is summarized in Table 1.
Instruction Formatting. After dataset aggregation, each sample is converted into an instruction-following format using a unified conversion scheme (see Section-C in the supplementary material). It should be noted that during dataset aggregation, some samples may share the same imagery from different datasets; however, since their textual instructions are regenerated under these rules, they are treated as distinct entries. In total, the resulting GeoZero-Raw dataset contains about 750K samples.
@@ 3.2.2 GeoZero-Instruct and GeoZero-Hard
We design a three-step process to build these datasets: (1) train the MLLM on GeoZero-Raw to obtain a data filtering model (DFM); (2) select difficult samples to construct GeoZero-Hard; and (3) perform data filtering based on GeoZero-Hard to produce GeoZero-Instruct. Intuitively, according to the maximum likelihood principle, samples mispredicted in a single inference pass are more likely to be difficult. Therefore, to ensure both the difficulty of GeoZero-Hard and the efficiency of sample selection, we adopt a coarse-to-fine two-stage filtering strategy. First, all samples in GeoZero-Raw are fed into the DFM for one inference pass, and only those incorrectly predicted are retained as potential hard-sample candidates. Next, these candidates are re-evaluated three times using the same model with different random seeds. Finally, we rank all candidates by their average error rate and select the top portion as the final hard samples. To maintain task balance, an equal number of samples are drawn from each of the four primary tasks. The selected samples collectively constitute the hard-sample pool, referred to as GeoZero-Hard. More details on the DFM and the hard-sample filtering procedure are provided in Section-D of the supplementary material.
Notably, to avoid potential image overlap within GeoZero-Raw (as discussed in Section-3.2.1), we remove all samples sharing the same image with GeoZero-Hard. The remaining samples form GeoZero-Instruct. As a result, GeoZero-Hard and GeoZero-Instruct contain approximately 20K and 610K samples, respectively.
@@ 3.3.1 Reward Function
Answer-Modulated Thinking Reward. AMTR begins with a thinking-quality score $s_{t}$, which comprehensively evaluates multiple aspects of the reasoning process, such as trajectory length, semantic diversity, and sentence redundancy. More details on $s_{t}$ are provided in the supplementary material (Section-F).
@@
Task-Specific Answer Reward. To adapt MLLMs to remote sensing vision-language tasks, following existing practices [27, 14, 17], we additionally design task-specific answer rewards $r_{a}$ for SC, VG, VQA, and IC. For consistency in reward computation, $r_{a}$ is expected to fall within the range of $[0,1]$. Moreover, $r_{a}$ should preferably be continuous so that it can smoothly reflect the quality differences among various answers, thereby facilitating model optimization during RL training. To this end, we formulate dedicated reward functions for each task, and the detailed formulations are provided in the supplementary material (Section-G).
@@ 3.3.2 Optimization Objective
In the original GRPO framework [18], the optimization objective depends on the logits of responses containing both the reasoning trajectory and the final answer. However, since our goal is to incentivize MLLMs to think from scratch, i.e., without exposure to any CoT data, the reference model (the MLLM after SFT) does not explicitly generate reasoning text. In this case, directly applying GRPO’s original formulation may inadvertently constrain the reasoning process. To address this issue, we refine the optimization objective in A2GRPO as follows:
@@
The definitions of $G$, $\varepsilon$, $\beta$, and the advantage term $A_{i}(r_{1},\dots,r_{G})$ follow the original GRPO [18], where $r_{i}$ denotes the overall reward (see Equation (3)) for each candidate response. The key difference lies in the introduction of a binary mask $M_{i}=\{m_{i,t}\}_{t=1}^{|o_{i}|}$, where $m_{i,t}=1$ if token $t$ is located within the answer tags (as detailed later), and $m_{i,t}=0$ otherwise. This design ensures that the distributional alignment with the reference model is enforced only on the answer portion, thereby preventing the reasoning tokens from being overly regularized and encouraging MLLMs to freely explore diverse reasoning trajectories.
@@ 3.3.3 Prompt template
Despite these efforts, MLLMs may still fail to generate explicit reasoning texts. In our experiments, we observed that directly using the original GRPO template [18] (<think>reasoning process</think><answer>answer</answer>) often fails to activate reasoning behaviors, as tokens ”<think>” or ”<” rarely appear at the beginning of sentences for SFT. Consequently, the model struggles to initiate coherent reasoning from scratch. To alleviate this issue, we simplify the GRPO prompt template by removing corresponding constraints while retaining only the ”<answer></answer>” markers. More details of our system prompts are provided in Section-E of the supplementary material.